HRMS, 28th Colloquium, Dijon, France, Aug. 28-Sep. 1, 2023

HIGH-RESOLUTION ROVIBRATIONAL SPECTROSCOPY OF CH^{35}Cl_{3}: CLUSTERING EFFECTS IN THE ν_{4} AND ν_{4} − ν_{3} BANDS

\underline{ADINA CEAUSU-VELCESCU} , Université de Perpignan, Laboratoire de Mathématiques et Physique, 52 Avenue Paul Alduy, 66860 Perpignan Cedex,

France; \underline{LAURENT MANCERON} , Ligne AILES, Synchrotron Soleil, L’Orme des Merisiers, St-Aubin, BP48, 91192 Gif-sur-Yvette, France; Université Paris-Cité and

Univ. Paris Est Créteil, CNRS, LISA, F-75013 Paris, France; HELMUT BECKERS, Institut für Chemie und Biochemie, Freie Universität Berlin, Fabeckstrasse 34-36, D-14195 Berlin, Germany; PIERRE GHESQUIÈRE, Manufacture Française des Pneumatiques Michelin, Centre de Technologie de Ladoux - Rue Orange - 63118 - Cébazat, France; PETR PRACNA, University of Chemistry and Technology, Department of Analytical Chemistry, Technická 5, 166

28 Prague 6, Czech Republic

The v_{4} = 1 fundamental vibration of CH^{35}Cl_{3} (C-H bending, E symmetry) has been revisited at high-resolution. For this purpose, three FTIR spectra, recorded respectively in the regions of the ν_{4} (1220 cm^{−1}), ν_{4} − ν_{3} (853 cm^{−1}), and ν_{3} (367 cm^{−1}) bands were employed. Spectra were all recorded at the Synchrotron Soleil, at a resolution of 0.001 cm^{−1} but with various optical pathlengths (12, 24, and 84 meters, respectively), using a monoisotopic CH^{35}Cl_{3} sample (99 % purity). As can be seen in the figure below, both the ν_{4} and ν_{4} − ν_{3} bands show remarkable clustering effects, due to the smallness of some combinations of rovibrational parameters.

p-number: p084  Submitted on Mon, 12 Jun 2023 15:34:02 +0200

These clustering effects are sometimes source of misassignments, especially when the systematic use of the lower state combination difference as a checking is not possible. In the present work, the assignment checking has been done through fundamental and difference transitions sharing a common upper level. This procedure allowed us to systematically correct and extend the K-assignments of the ^{r}R_{K}(J) transitions in the ν_{4} band, with respect to the previous high-resolution study^{1}. The least-squares fit, performed on more than 6900 transitions of the ν_{4} and ν_{4} −ν_{3} bands (0 ≤ J ≤ 100 and −75 ≤ K · ΔK ≤ 75), combined with 800 IR (ν_{3}) and 1300 MMW (v_{3} = 1) data, allowed to obtain an accurate description of the v_{4} = 1 level (global standard deviation of 0.160 cm^{−1}).

1R. Anttila, S. Alanko, V.-M. Horneman, Mol. Phys. 102 (2004) 1537-1542.

p-number: p084  Submitted on Mon, 12 Jun 2023 15:34:02 +0200